# Algebraic non-Hermitian skin effect and unified non-Bloch band theory in arbitrary dimensions

### v1

@@ III.3 The exact 2D open-boundary wavefunction from GFS basis
Here, we show that the analytic wavefunction, constructed from the GFS according to our formula, fully agrees with the numerical wavefunction obtained from the direct diagonalization of the Hamiltonian matrix, as shown in Figs. 3(d) and (d1)-(d3), thereby validating our 2D GBZ formula.
@@
In the model Hamiltonian given by Eq. (21), the GFS corresponding to $E_{0}$ is obtained and shown in Figs. 3(b) and (c). For a given $\rho_{y}$, there exist two non-Bloch waves, $\beta_{x,1}(\rho_{y})$ and $\beta_{x,2}(\rho_{y})$, that have the same amplitude $|\beta_{x,1}(\rho_{y})|=|\beta_{x,2}(\rho_{y})|$. These non-Bloch waves combine to form a standing wave that satisfies OBCs in the $x$ direction, which is expressed as $\phi_{\rho_{y}}(x)=\beta^{x}_{x,1}(\rho_{y})-\beta^{x}_{x,2}(\rho_{y})$. An OBC wavefunction is the superposition of $2L_{x}$ such standing waves, each coupled with the associated $\rho_{y}$. Therefore, for this model, the standing-wave construction in Eq. 28 can be explicitly expressed as
$\Psi_{E_{0}}(\bm{r})=\sum\nolimits_{i=1}^{2L_{x}}A_{i}\varphi_{i}(x,y)=\sum\nolimits_{i=1}^{2L_{x}}A_{i}\,\rho_{i}^{y}\,[\beta^{x}_{i,1}-\beta^{x}_{i,2}],$ (31)
@@
where the subscripts $x$ and $y$ for $\beta$ and $\rho$ are omitted for clarity, and $\varphi_{i}(x,y)$ signifies the $i$-th component of the open-boundary wavefunction. The $2L_{x}$ pairs of non-Bloch wave components, $\left(\rho_{i};\beta_{i,1},\beta_{i,2}\right)|_{i=1,\dots,2L_{x}}$, constitute the GFS for $E_{0}$ and are represented by the red and blue dots in Figs. 3(b) and (c). The $2L_{x}$ superposition coefficients $\{A_{i=1,\dots,2L_{x}}\}$ are exactly determined by the $2L_{x}$ open-boundary conditions along the $y$ direction.
The open boundary conditions in the $y$ direction are obtained from the recursion equation Eq. (23) when $y=1$ and $y=L_{y}$. Plugging in $\psi_{0}=\psi_{L_{y}+1}=0$, we get
$\begin{split}(\textbf{h}_{0}-E_{0}\,\mathbb{I}_{L_{x}})\,\psi_{1}+\textbf{h}_{1}\,\psi_{2}&=0;\\ \textbf{h}_{-1}\,\psi_{L_{y}-1}+(\textbf{h}_{0}-E_{0}\,\mathbb{I}_{L_{x}})\,\psi_{L_{y}}&=0.\end{split}$ (32)
Here, the layer wavefunction is expressed as $\psi_{y}=\{\Psi_{E_{0}}(x{=}1,y),\dots,\ \Psi_{E_{0}}(x{=}L_{x},y)\}^{T}$, with each component $\Psi_{E_{0}}(x,y)$ given by Eq. (31). Therefore, these two boundary equations in Eq. (32) encompass $2L_{x}$ constraints on the wavefunction components. These constraints can be conveniently expressed in matrix form as follows:
$\mathcal{M}_{B}({E_{0}})\,\textbf{A}=0,$ (33)
where A is a column vector of superposition coefficients $\{A_{1},\dots,A_{i},\dots,A_{2L_{x}}\}$, and $\mathcal{M}_{B}$ is a $2L_{x}\times 2L_{x}$ boundary matrix.
Here, the boundary matrix provides a criterion that the determinant of the boundary matrix vanishes,
$\det[\mathcal{M}_{B}(E_{0})]=0,$ (34)
if and only if ${E_{0}}$ belongs to open-boundary eigenvalues. Correspondingly, the zero-energy (null) eigenvector A of the boundary matrix $\mathcal{M}_{B}({E_{0}})$ precisely determines superposition coefficients $\{A_{i}\}$. As an illustration, we calculate these superposition coefficients using the boundary matrix for energy $E_{0}=0.819-1.108i$, and obtain the constructed analytical OBC wavefunction, as shown in Fig. 3(d). As indicated by the selected three generic layers in Fig. 3(d), we compare the constructed analytical wavefunction (the blue lines) with the numerical wavefunction (the red dots) that is computed by diagonalizing the Hamiltonian matrix. The exact match confirms the validity of our established GBZ formula.
III.4 The non-uniqueness of GBZ in higher dimensions
Unlike in one dimension, the GBZ (or GFS) in higher dimensions is not unique and depends on the choice of the layer basis for a given Hamiltonian. From the perspective of the wavefunction, in one dimension, the GBZ condition exactly provides the standing wave condition: non-Bloch waves that satisfy this GBZ condition can form stable standing waves under open-boundary conditions. However, in two dimensions, dimension analysis shows that a 3D GBZ manifold needs only one standing wave condition. This means that the wavefunction can only form standing waves along either one spatial dimension, instead of in both directions. Consequently, the choices of standing wave bases in different directions correspond to distinct GBZ manifolds.
@@ III.4 The non-uniqueness of GBZ in higher dimensions
For example, the Hamiltonian in Eq. (21) can also be expressed within the layer-$x$ basis. Following a similar procedure, the non-Bloch layer transfer matrix along the $x$ direction $\mathbb{T}(\beta_{y},E)$ and the bulk equation $f(\rho_{x},\beta_{y},E)=0$ can be derived. Here, $\rho_{x}$ indicates the eigenvalues of the transfer matrix along the $x$ direction. For the example in Eq. (21), the corresponding GBZ condition under layer-$x$ basis becomes:
@@
Compared it with the GBZ condition given by Eq. (30) under the layer-$y$ basis, two different GBZ conditions lead to two distinct GBZs. In the layer-$x$ basis, the projection of corresponding GFS onto $\rho_{x}$ plane forms arcs while the projection onto $\beta_{y}$ plane forms closed curves that satisfy the GBZ condition in Eq. (35). Using this basis, the open-boundary wavefunction can be constructed by a superposition of standing waves in the $y$ direction, each coupled with the corresponding transfer matrix eigenvalue $\rho_{x}$
$\begin{split}\Psi_{E}(x,y)=\sum\nolimits_{\rho_{x}\in\text{GFS}}A(\rho_{x})\,\rho^{x}_{x}\,\phi_{\rho_{x}}(y),\end{split}$ (36)
where $\phi_{\rho_{x}}(y)$ indicates the standing wave coupled with the transfer matrix eigenvalue $\rho_{x}$. The superimposition coefficients of these components are determined by the open-boundary conditions in the $x$ direction. It is important to note that though the GBZ bases obtained from the layer-$x$ and layer-$y$ transfer matrix are different, the superposition coefficients are adjusted accordingly. This ensures that the same OBC wavefunction $\Psi_{E}(x,y)$ and other physical quantities are ultimately produced.
@@ III.5 Discussion on other scenarios in two dimensions
Here, we discuss other 2D scenarios that require fine-tuning or occupy a parameter space of zero measure. For completeness, all 2D cases can be classified based on the coverage of the OBC spectrum and the dimensionality of GBZ. There are four potential scenarios for their combinations: $\dim{\sigma_{\text{OBC}}}=2$ or $1$ and $\dim{\text{GBZ}}=3$ or $2$. Previously, we demonstrated that the most generic 2D scenario features a 3D GBZ manifold and area-form OBC spectrum. Other cases, either disallowed or necessitating fine-tuning, are discussed in this section.
@@
When the OBC eigenvalues span a finite area and $\dim{\text{GBZ}}=2$ (Fig. 1), the dimension of GFS reduces to 0, based on the relation in Eq. (5). The zero-dimensional GFS means that OBC eigenstates are formed by a discrete number of Bloch or non-Bloch waves and thus manifest as extended bulk waves or exponentially localized skin modes. This generally requires that the Hamiltonian respects certain spatial symmetry and that the open boundary geometry matches this symmetry. A representative example is as follows. For the example in Eq. (21), we set the parameters $t_{xy}=0,t_{x}=1$, and $t_{y}=i$, such that the Hamiltonian is decoupled in the $x$ and $y$ directions
$\mathcal{H}(\beta_{x},\beta_{y})=\beta_{x}+\beta_{x}^{-1}+i(\beta_{y}+\beta_{y}^{-1}).$ (37)
@@
For a specific complex energy, say $E_{0}=1+i$, the $\beta_{x}$ and $\beta_{y}$ can be solved as four points where $|\beta_{x}|=|\beta_{y}|=1$. In this case, GFS is composed of these four Bloch points, thus the GFS reduces from the 1D curves to 0D points. Correspondingly, the GBZ conditions have two constraints, $|\beta_{x,1}(\beta_{y},E)|=|\beta_{x,2}(\beta_{y},E)|$ and $|\beta_{y,1}(\beta_{x},E)|=|\beta_{y,2}(\beta_{x},E)|$, and the standing waves can be enforced simultaneously along both the $x$ and $y$ directions. Since the GFS is 0D, when the energy scans through the entire continuum spectrum, the corresponding GBZ is a 2D manifold.
@@
When the OBC eigenvalues form some arcs within the complex energy plane, the GBZ manifests as a 2D manifold. This implies that standing wave conditions can be enforced in both the $x$ and $y$ directions. We provide an intuitive argument by contradiction as follows. We begin with the assumption that $\dim{\text{GBZ}}=3$ in this arc-spectrum case. According to the relation in Eq. (5), GFS is a 2D manifold, i.e., $\dim{\text{GFS}}=2$. However, a 3D GBZ manifold would suggest that only one standing wave condition can be enforced based on Eq. (5). Under these conditions, once the layer-$y$ basis is selected, the OBC eigen wavefunctions are constructed based on Eq. (28). Contrasting with our initial supposition, nevertheless, the wavefunction construction here only requires order-$L$ non-Bloch solutions on the GFS, suggesting that the actual GFS consists of 1D curves, i.e., $\dim{\text{GFS}}=1$. Consequently, we arrive at the contradiction with the initial assumption that $\dim{\text{GBZ}}=3$ in the arc-spectrum case. Therefore, we conclude that in arc-spectrum scenarios, the GBZ manifold is 2-dimensional. One example is the 2D Hatano-Nelson model, which is given by
@@
with the real valued $t_{i},w_{i}>0$ and $t_{i}\neq w_{i}$ for $i=x,y$. The GBZ can be directly solved through an imaginary gauge transformation: $\beta_{x}\to\sqrt{w_{x}/t_{x}}\,\widetilde{\beta}_{x}$, and $\beta_{y}\to\sqrt{w_{y}/t_{y}}\,\widetilde{\beta}_{y}$. After this transformation, the 2D non-Hermitian Hamiltonian is mapped to a Hermitian Hamiltonian in the basis of $\widetilde{\beta}_{x}$ and $\widetilde{\beta}_{y}$: $\mathcal{H}(\widetilde{\beta}_{x},\widetilde{\beta}_{y})=\sqrt{t_{x}w_{x}}(\widetilde{\beta}_{x}+\widetilde{\beta}_{x}^{-1})+\sqrt{t_{y}w_{y}}(\widetilde{\beta}_{y}+\widetilde{\beta}_{y}^{-1})$. For the Hermitian Hamiltonian $\mathcal{H}(\widetilde{\beta}_{x},\widetilde{\beta}_{y})$, the GBZ coincides with the BZ, and its GFS reduces to a $(d-1)$-dimensional Fermi surface. The OBC eigen wavefunctions are Bloch wave standing waves along both $x$ and $y$ directions. Upon transforming this Hermitian Hamiltonian back to the non-Hermitian one, we can obtain the GBZ in the original non-Hermitian Hamiltonian is a 2D distorted torus with radius $|\beta_{i}|=\sqrt{w_{i}/t_{i}}$ for $i=x,y$. In this case, the OBC eigenstates are Bloch waves modified by an exponential prefactor, causing the wavefunctions to concentrate at the corners of open boundaries.
So far, we have exhausted all conceivable scenarios in two dimensions. It’s worth noting that in these two specific fine-tuning scenarios discussed, both the GBZ and GFS can be solved through the layer transfer matrix approach. However, in these specific scenarios, the GBZ and GFS are independent of the choice of layer basis and the GBZ manifold is unique, contrasting with the non-uniqueness of GBZ for generic 2D systems. The uniqueness of the GBZ manifold here stems from the alignment of the number of standing wave conditions with the spatial dimensions. Given that these scenarios require precise fine-tuning, our primary findings in this study are focused on the more general two-dimensional systems.

### v2

@@ III.3 The 2D wavefunction construction from GFS bases under open boundary conditions
Here, we demonstrate that the ansatz wavefunction constructed from the GFS basis fully agrees with the result obtained from direct numerical diagonalization of the $L^{2}\times L^{2}$ Hamiltonian matrix, as shown in Fig. 4(d), thereby validating our 2D GFS formula.
Notably, the GFS framework classifies the $d$-dimensional systems into $d$ distinct cases based on the standing-wave construction, as illustrated in Fig. 1. In our model given by Eq. (21), the standing waves can be enforced along either the $x$ or $y$ direction, placing it into $n=1$ case, where $n$ denotes the number of standing wave conditions. Our framework effectively reduces the complexity of an order-$L^{2}$ diagonalization (solving an $L^{2}\times L^{2}$ Hamiltonian matrix derived from the original Schrödinger eigenequation) to the complexity of an order-$L$ diagonalization (solving an $L\times L$ sub-extensive boundary matrix). More generally, the system can be reduced to an order-$L^{d-n}$ diagonalization problem. When $n=d$, order-$L^{0}$ boundary matrix is needed, indicating that in this case boundary details are irrelevant in the thermodynamic limit. The 1D non-Bloch band theory falls into this category, where $n=d=1$. Otherwise, when $n<d$, solving an order-$L^{d-n}$ boundary matrix becomes necessary, which underscores the crucial role of boundary conditions in these scenarios.
@@
In the model Hamiltonian given by Eq. (21), the GFS corresponding to $E_{0}$ is obtained and its projections onto complex $\beta_{x}$ and $\rho_{y}$ planes are shown in Figs. 4(b) and (c). For a given $\rho_{y}$, there exist two non-Bloch waves, $\beta_{x,1}(\rho_{y})$ and $\beta_{x,2}(\rho_{y})$, that have the same amplitude $|\beta_{x,1}(\rho_{y})|=|\beta_{x,2}(\rho_{y})|$. These non-Bloch waves combine to form standing waves that satisfy OBCs in the $x$ direction, expressed as $\varphi_{\rho_{y}}(x)=\beta^{x}_{x,1}(\rho_{y})+B\beta^{x}_{x,2}(\rho_{y})$, where $B$ is a superposition coefficient. The OBCs in the $x$ direction impose the conditions $\varphi_{\rho_{y}}(x=0)=\varphi_{\rho_{y}}(x=L_{x}+1)=0$ on the standing-wave component. Specifically, $\varphi_{\rho_{y}}(x=0)=0$ leads to B=1B=-1, and $\varphi_{\rho_{y}}(x=L_{x}+1)=0$ discretizes the continuous GFS curve into discrete points, with the number of points matching the boundary degrees of freedom. See more details in Appendix A. An OBC wavefunction is the superposition of $2L_{x}$ such standing wave components, each coupled with the associated $\rho_{y}(\beta_{x},E_{0})$. Therefore, for this model, the standing-wave construction in Eq. (28) can be explicitly expressed as
$\Psi_{E_{0}}(\bm{r})=\sum\nolimits_{i=1}^{2L_{x}}A_{i}\,\rho_{i}^{y}\,\varphi_{i}(x)=\sum\nolimits_{i=1}^{2L_{x}}A_{i}\,\rho_{i}^{y}\,[\beta^{x}_{i,1}-\beta^{x}_{i,2}],$ (31)
where the subscripts $x$ and $y$ for $\beta$ and $\rho$ are omitted for simplicity, and $\varphi_{i}(x)$ signifies the $i$-th standing-wave component associated with $\rho_{i}$.
@@
The $2L_{x}$ pairs of non-Bloch wave components, $\left(\rho_{i};\beta_{i,1},\beta_{i,2}\right)|_{i=1,\dots,2L_{x}}$, constitute the discretized GFS for $E_{0}$ and are represented by the red and blue dots in Figs. 4(b) and (c). The $2L_{x}$ superposition coefficients $\{A_{i=1,\dots,2L_{x}}\}$ are exactly determined by the $2L_{x}$ open-boundary conditions in the $y$ direction. The open boundary conditions in the $y$ direction are defined by the specific configurations on the boundary layers at $y=1$ and $y=L_{y}$. More generally, these open boundary conditions can be encapsulated in the following equations:
$\begin{split}(\textbf{b}_{1}-E_{0}\,\mathbb{I}_{L_{x}})\,\psi_{1}+\textbf{b}_{+}\,\psi_{2}&=\textbf{0};\\ \textbf{b}_{-}\,\psi_{L_{y}-1}+(\textbf{b}_{L_{y}}-E_{0}\,\mathbb{I}_{L_{x}})\,\psi_{L_{y}}&=\textbf{0}.\end{split}$ (32)
Here, $\psi_{i}$ represents a column of wavefunction components at layer $y=i$, each of which includes $L_{x}$ components. Specifically, $\textbf{b}_{1}$ and $\textbf{b}_{L_{y}}$ are the $L_{x}\times L_{x}$ matrices, which are given by the specific boundary conditions at the layers $y=1$ and $y=L_{y}$. The $\textbf{b}_{+}$ ($\textbf{b}_{-}$) represents the adjacent matrix between boundary layer $y=1$ ($y=L_{y}$) and the bulk. In general, we can add the boundary disorders to modify the boundary conditions $\{\textbf{b}_{1},\textbf{b}_{L_{x}},\textbf{b}_{\pm}\}$. The case with boundary disorders is discussed in Section VI. For notational clarity, we label the standing-wave component as
$\varphi_{x,j}\equiv\beta^{x}_{1}(\rho_{j})-\beta^{x}_{2}(\rho_{j}),$
[…]
$\begin{split}&\mathcal{B}_{ij}=\sum_{x}\{[(\textbf{b}_{1})_{i,x}-E_{0}\delta_{i,x}]\rho_{j}+(\textbf{b}_{+})_{i,x}\rho_{j}^{2}\}\,\varphi_{x,j};\\ &\mathcal{C}_{ij}=\sum_{x}\{(\textbf{b}_{-})_{i,x}\rho_{j}^{L_{y}-1}+[(\text{b}_{L_{y}})_{i,x}-E_{0}\delta_{i,x}]\rho^{L_{y}}_{j}\}\,\varphi_{x,j}.\end{split}$
The subscripts $i,x$ corresponds to the lattice sites and runs from $1$ to $L_{x}$, and the index $j$ of $\rho_{j}$ corresponds to discretized GFS components and runs from $1$ to $2L_{x}$. Therefore, $\mathcal{B}$ and $\mathcal{C}$ are two rectangular matrices with size of $L_{x}\times 2L_{x}$. With these notations, the boundary conditions in Eq. (32) can be expressed in a matrix form:
$\mathcal{M}_{B}(E_{0})\,\textbf{A}=\begin{pmatrix}\mathcal{B}_{1,1}&\mathcal{B}_{1,2}&\dots&\mathcal{B}_{1,L_{x}}&\mathcal{B}_{1,L_{x}+1}&\dots&\dots&\mathcal{B}_{1,2L_{x}}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ \mathcal{B}_{L_{x},1}&\mathcal{B}_{L_{x},2}&\dots&\mathcal{B}_{L_{x},L_{x}}&\mathcal{B}_{L_{x},L_{x}+1}&\dots&\dots&\mathcal{B}_{L_{x},2L_{x}}\\ \mathcal{C}_{1,1}&\mathcal{C}_{1,2}&\dots&\mathcal{C}_{1,L_{x}}&\mathcal{C}_{1,L_{x}+1}&\dots&\dots&\mathcal{C}_{1,2L_{x}}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ \mathcal{C}_{L_{x},1}&\mathcal{C}_{L_{x},2}&\dots&\mathcal{C}_{L_{x},L_{x}}&\mathcal{C}_{L_{x},L_{x}+1}&\dots&\dots&\mathcal{C}_{L_{x},2L_{x}}\\ \end{pmatrix}\begin{pmatrix}A_{1}\\ \vdots\\ A_{L_{x}}\\ A_{L_{x}+1}\\ \vdots\\ A_{2L_{x}}\end{pmatrix}=\textbf{0}.$ (33)
The boundary matrix $\mathcal{M}_{B}(E_{0})$ is a $2L_{x}\times 2L_{x}$ matrix that depends on the energy $E_{0}$. For the boundary matrix $\mathcal{M}_{B}$, the first $L_{x}$ rows correspond to the boundary conditions at $y=1$ (the lower boundary), while the last $L_{x}$ rows correspond to the boundary conditions at $y=L_{y}$ (the upper boundary). The $2L_{x}$ columns of the boundary matrix correspond to the $2L_{x}$ discretized GFS components. The null space of the boundary matrix $\mathcal{M}_{B}(E_{0})$ determines the superposition coefficients $\textbf{A}=\{A_{1},\dots,A_{i},\dots,A_{2L_{x}}\}^{T}$ in Eq. (31). With this, the ansatz wavefunction constructed from the GFS basis is fully determined. Additionally, the boundary matrix provides a criterion for determining open-boundary eigenvalues:
$\det[\mathcal{M}_{B}(E_{0})]=0$ (34)
if and only if $E_{0}$} belongs to open-boundary spectrum.
As an illustration, we compute the superposition coefficients using the boundary matrix at energy $E_{0}=0.819-1.108i$, and construct the ansatz OBC wavefunction, as shown in Fig. 4(d). This wavefunction is analytical in the sense that its expression in Eq. (31) is derived from the GFS basis, with the corresponding GFS curve obtained analytically via the resultant method (see Appendix A for details). Thus, the constructed wavefunction is fundamentally distinct from the numerical wavefunction obtained by direct diagonalization of the full Hamiltonian matrix. To validate our approach, we compare the ansatz wavefunction (blue lines) with the numerically computed wavefunction (red dots) at three representative layers in Fig. 4(d). The perfect agreement between the two confirms the accuracy of the GFS-based formulation.
III.4 The non-uniqueness of GFS in higher dimensions
Unlike in one dimension, the GFS in higher-dimensional systems may not be uniquely defined for a given Hamiltonian. This non-uniqueness arises from the wide variety of possible boundary conditions or geometries in higher dimensions, for instance, the infinite family of parallelogram-shaped boundaries in 2D. When a system is sensitive to boundary geometry, as reflected in its spectrum or wavefunction, a GFS theory must explicitly account for these boundary-specific features. As a result, the GFS becomes inherently non-unique. However, for certain special systems that can be decomposed into independent 1D subsystems, the GFS or GBZ can be constructed as a direct product of 1D solutions, rendering them uniquely defined in such cases.
We now provide an intuitive explanation for this non-uniqueness based on the standing-wave conditions. In a $d$-dimensional system, if the number of imposed standing-wave conditions (denoted by $n$) is less than the spatial dimension $d$, i.e., $n<d$, the GBZ and GFS manifolds are not uniquely determined by the bulk Hamiltonian. This ambiguity arises from the freedom in choosing the directions along which standing waves are imposed. For instance, in two dimensions with $n=1$, one may impose the standing-wave condition along either the $x$- or $y$-direction. Each choice yields a complete non-Bloch basis capable of reconstructing the OBC wavefunction, but the resulting GBZ and GFS manifolds are generally incompatible and distinct. An explicit example illustrating this non-uniqueness is provided in Appendix B. In contrast, when $n=d$, the standing-wave conditions fully constrain the system, leaving only one consistent way to construct the non-Bloch basis. As a result, the GBZ and GFS manifolds are uniquely defined.
@@ III.4 The non-uniqueness of GFS in higher dimensions
As an example, the Hamiltonian in Eq. (21) can also be expressed in the layer-$x$ basis. Following a similar procedure, one can derive the non-Bloch layer transfer matrix along the $x$-direction, denoted by $\mathbb{T}(\beta_{y},E)$, along with the corresponding bulk equation $f(\rho_{x},\beta_{y},E)=0$, where $\rho_{x}$ represents the eigenvalues of the transfer matrix along $x$ direction. For this representation, the GFS condition becomes
@@
Compared to the GFS condition in Eq. (30) under the layer-$y$ basis, this alternative leads to a distinct GFS curve. In the layer-x basis, the projection of the GFS onto the $\rho_{x}$-plane forms arcs, while its projection onto the $\beta_{y}$-plane yields closed curves that satisfy the condition in Eq. (35). Within this basis, the open-boundary wavefunction can be constructed as a superposition of standing waves along the y-direction, each coupled to a corresponding transfer eigenvalue $\rho_{x}$:
$\begin{split}\Psi_{E}(x,y)=\sum\nolimits_{\rho_{x}\in\text{GFS}}A(\rho_{x})\,\rho^{x}_{x}\,\varphi_{\rho_{x}}(y),\end{split}$ (36)
where $\varphi_{\rho_{x}}(y)$ indicates the standing wave coupled with the transferring component $\rho_{x}$. The coefficients $A(\rho_{x})$ are determined by the open-boundary conditions along the $x$-direction. It is important to emphasize that although the GFS bases obtained via standing-wave conditions along the $x$- and $y$-directions differ, they ultimately yield the same OBC eigenvalues and eigenstates.
@@ III.5 Discussion on other scenarios in two dimensions
Here, we consider special 2D cases involving symmetry-protected bulk Hamiltonians or particular boundary geometries, which require fine-tuning and occupy a Hamiltonian parameter space of measure zero. Based on the dimensionality of the GBZ and the coverage of the OBC spectrum, there are four possible scenarios: $\dim{\sigma_{\text{OBC}}}=2$ or $1$ and $\dim{\text{GBZ}}=3$ or $2$. The most generic 2D case, as previously shown, features a 3D GBZ manifold and a 2D (area-covering) OBC spectrum. The remaining cases, which are either forbidden by general constraints or only realized through fine-tuning, will be discussed in this section.
@@
We first consider the case where the OBC eigenvalues span a finite area (i.e., $\dim\sigma_{\text{OBC}}=2$) and the GBZ is two-dimensional (i.e., $\dim\text{GBZ}=2$), as illustrated in Fig. 2. According to the dimensional relation in Eq. (5), the corresponding GFS in this case reduces to a set of isolated points, i.e., $\dim\text{GFS}=0$. The zero-dimensional GFS means that OBC eigenstates are formed by a discrete number of Bloch or non-Bloch waves and thus manifest as extended bulk waves or exponentially localized skin modes. This generally requires that the Hamiltonian respects certain spatial symmetry and that the open boundary geometry matches this symmetry. A representative example is as follows. For the example in Eq. (21), we set the parameters $t_{xy}=0,t_{x}=1$, and $t_{y}=i$, such that the Hamiltonian is decoupled in the $x$ and $y$ directions
$\mathcal{H}(\beta_{x},\beta_{y})=\beta_{x}+\beta_{x}^{-1}+i(\beta_{y}+\beta_{y}^{-1})$ (37)
@@
with open boundary along the $x$ and $y$ directions, i.e., square boundary geometry. For a fixed complex energy, say $E_{0}=1+i$, the $\beta_{x}$ and $\beta_{y}$ can be solved as four points where $|\beta_{x}|=|\beta_{y}|=1$. In this case, the GFS consists of these four isolated Bloch points, reducing from 1D curves to 0D points. Correspondingly, the GFS conditions include two constraints, $|\beta_{x,1}(\beta_{y},E)|=|\beta_{x,2}(\beta_{y},E)|$ and $|\beta_{y,1}(\beta_{x},E)|=|\beta_{y,2}(\beta_{x},E)|$, indicating that standing-wave conditions can be simultaneously enforced along both the $x$- and $y$-directions. Since the GFS is 0D for each fixed energy, scanning over the full energy continuum generates a 2D GBZ manifold.
@@
The case with $\dim\sigma_{\text{OBC}}=1$ and $\dim{\text{GBZ}}=3$ is impossible for the following reason. According to the relation in Eq. (5), GFS is a 2D manifold, i.e., $\dim{\text{GFS}}=2$ in this case. However, a 3D GBZ manifold would suggest that only one standing wave condition can be enforced based on Eq. (5). Under these conditions, once the layer-$y$ basis is selected, the OBC eigenstates are constructed based on Eq. (28). While the wavefunction construction here only requires non-Bloch solutions with the number of order-$L$ on the GFS, suggesting that the actual GFS is 1D curves, i.e., $\dim{\text{GFS}}=1$. Consequently, there are contradictions with $\dim{\text{GFS}}=2$ the case suggests, making this case impossible.
The remaining possible case is $\dim\sigma_{\text{OBC}}=1$ and $\dim{\text{GBZ}}=2$. One representative example is the 2D Hatano-Nelson model, whose Hamiltonian is given by
@@
with the real valued $t_{i},w_{i}>0$ and $t_{i}\neq w_{i}$ for $i=x,y$. The GBZ can be directly solved through an imaginary gauge transformation: $\beta_{x}\to\sqrt{w_{x}/t_{x}}\,\widetilde{\beta}_{x}$, and $\beta_{y}\to\sqrt{w_{y}/t_{y}}\,\widetilde{\beta}_{y}$. After this transformation, the 2D non-Hermitian Hamiltonian is mapped to a Hermitian Hamiltonian in the basis of $\widetilde{\beta}_{x}$ and $\widetilde{\beta}_{y}$: $\mathcal{H}(\widetilde{\beta}_{x},\widetilde{\beta}_{y})=\sqrt{t_{x}w_{x}}(\widetilde{\beta}_{x}+\widetilde{\beta}_{x}^{-1})+\sqrt{t_{y}w_{y}}(\widetilde{\beta}_{y}+\widetilde{\beta}_{y}^{-1})$. For the Hermitian Hamiltonian $\mathcal{H}(\widetilde{\beta}_{x},\widetilde{\beta}_{y})$, the GBZ coincides with the BZ, and its GFS reduces to a 1D Fermi surface. The OBC eigenstates are Bloch wave standing waves along both $x$ and $y$ directions. Upon transforming this Hermitian Hamiltonian back to the non-Hermitian one, we can obtain the GBZ in the original non-Hermitian Hamiltonian as a 2D distorted torus with radius $|\beta_{i}|=\sqrt{w_{i}/t_{i}}$ for $i=x,y$. In this case, the OBC eigenstates are Bloch waves modified by an exponential prefactor, causing the wavefunctions to concentrate at the corners of open boundaries.
We have now discussed other possible scenarios in two dimensions, all of which are compatible with and can be described by the established GFS framework. Notably, in the two fine-tuned cases considered above, both the GBZ and GFS are uniquely determined, as the standing-wave conditions can be simultaneously imposed along both spatial directions, eliminating any redundancy in the GFS basis. In these cases, the OBC wavefunctions in both cases exhibit either extended Bloch waves or exponentially localized skin modes, due to the 0-dimensional GFS ($\dim\text{GFS}=0$), as shown in Fig. 2.